ADNAN MUHAMMED

Grade 12,

A standing wave y=Asin(20\3¶x)cos(1000¶t) is maintained in a taut string where y and x are expressed in metres. The distance between the successive points oscillating with the amplitude A\2 across a node is equal to

A standing wave y=Asin(20\3¶x)cos(1000¶t) is maintained in a taut string where y and x are expressed in metres. The distance between the successive points oscillating with the amplitude A\2 across a node is equal to

Grade:12

1 Answers

Suresh Kumar Chhapola
6 years ago
To find req. Distance we have to find the distance of point have amplitude A/2 from the node and double of it will be required distanceAsin(--)=A/220/3pix=pi/6x=2.5cmRequired distance=2x=2*2.5=5cm